# Improving MLLMs in Embodied Exploration and Question Answering with Human-Inspired Memory Modeling

### v1

@@ Abstract
Deploying Multimodal Large Language Models as the brain of embodied agents remains challenging, particularly under long-horizon observations and limited context budgets. Existing memory-assisted methods often rely on textual summaries, which discard rich visual and spatial details and remain brittle in non-stationary environments. In this work, we propose a non-parametric memory framework that explicitly disentangles episodic and semantic memory for embodied exploration and question answering. Our retrieval-first, reasoning-assisted paradigm recalls episodic experiences via semantic similarity and verifies them through visual reasoning, enabling robust reuse of past observations without rigid geometric alignment. In parallel, we introduce a program-style rule extraction mechanism that converts experiences into structured, reusable semantic memory, facilitating cross-environment generalization. Extensive experiments demonstrate state-of-the-art performance on embodied question answering and exploration benchmarks, yielding a 7.3% gain in LLM-Match and an 11.4% gain in LLM-MatchSPL on A-EQA, as well as +7.7% success rate and +6.8% SPL on GOAT-Bench. Analyses reveal that our episodic memory primarily improves exploration efficiency, while semantic memory strengthens complex reasoning of embodied agents.
@@ I Introduction
In recent years, Multimodal Large Language Models (MLLMs) have advanced rapidly, demonstrating strong reasoning and perception capabilities across a wide range of domains. Despite this progress, deploying MLLMs as the core decision-making module of embodied agents in real-world environments remains highly challenging. A key difficulty lies in enabling agents to efficiently process long-horizon observations while retaining task-relevant information under limited context budgets. To address this issue, recent works introduce various forms working memory, allowing agents to retrieve a small subset of relevant past observations instead of conditioning on the entire sensory history. By coupling task-relevant information retrieval with frontier-based exploration strategies, these methods achieve efficient and effective performance in embodied exploration and embodied question answering tasks.
@@
However, most existing approaches remain confined to information within the current episode, limiting their ability to leverage accumulated experience over time. In contrast, real-world embodied agents are expected to continuously integrate observations across episodes and exploit long-term knowledge acquired from past interactions. While [6] incorporates cross-episode memory, its task definition remain restrictive, typically assessing an MLLM-based information-collection system within a fixed and fully explored space derived from previous timeline. Moreover, approaches such as [30] and [31] fundamentally summarize cross-episode information into hierarchical textual descriptions, which are then used to guide future exploration. This design inevitably discards a substantial amount of spatial information and rich visual cues in raw RGB observations, constraining the agent’s ability to remember.
@@
To eflects real-world embodied scenario, we formulate the question in assumption that environments are non-stationary, and then reveals two fundamental challenges. First, maintaining long-term maps—whether represented as persistent nodes [25] or voxel grids [22]-becomes inherently difficult in the presence of environmental changes, sensor noise, and long-horizon drift. Second, information retrieved from past episodes does not necessarily reflect the current state of the environment. Based on these observations, we argue that an effective embodied agent should neither rely on rigid long-term geometric fusion nor blindly trust past episodic information. Instead, it should recall prior experiences in a soft and associative manner, while continuously grounding them in active exploration and perception — closely mirroring how humans reason about and interact with dynamic environments.
Motivated by this insight, we draw inspiration from human memory systems and propose a non-parametric memory framework that explicitly disentangles episodic and semantic memory for embodied exploration and question answering. First, our approach adopts retrieval-first and reasoning-assisted paradigm, where episodic memories are recalled based on semantic similarity and selectively verified through a system-2 visual reasoning. Second, rather than storing memory as free-form textual summaries which are often verbose, environment-specific, and difficult to reuse, we propose program-style enhanced rule extraction to form semantic memory. By generating pseudocode-style workflows and incorporating decision deviation detection, our approach extracts structured, high-level, and reusable decision rules from embodied experiences. These rules formed semantic memory generalize more effectively across environments, providing stronger and more reliable reasoning support during decision making.
We achieve new state-of-the-art results on embodied question answering and exploration, with 65.6% LLM-Match and 48.7% LLM-MatchSPL on A-EQA [15], and a 72.8% success rate with 56.1% SPL on GOAT-Bench [10]. Extensive experiments show that our proposed episodic memory improves exploration efficiency, while semantic memory enhances performance on complex, long-horizon reasoning tasks.
@@ II-A MLLM-based Embodied Agents
Despite these advances, foundation model still incur substantial computational costs, and naively processing uniformly sampled frames is often suboptimal for questions that require selectively identifying high-value observations for exploration and answering. To address these challenges, [7, 23, 25, 33] introduces CLIP and SAM-based open-vocabulary scene graphs to structure semantics for MLLM-driven exploration. While [27] further incorporates raw images to enrich memory, its key-frame selection remains constrained by detectors and lacks explicit geometric modeling, leading to inefficient long-horizon exploration. In contrast to existing geometric-semantic graph methods often depend on heavy multi-frame fusion, our approach marks a shift toward decoupling semantic and geometric spaces, enabling retrieval-time interaction that supports zero-shot generalization and adaptive cross-episode recalling.
@@ II-B Experience and Memory for Embodied Agents
Research on memory mechanisms for embodied agents seeks to enable the storage, retention, and retrieval of past experiences, facilitating the transition from purely reactive systems to agents capable of maintaining context and autonomous adaptation. [1] and [30] adopt description text, temporal and spatial cues to manage continuously growing histories, allowing robots to recall previously visited regions and reduce redundant exploration. However, memory built primarily on video or image captioning discard a substantial amount of fine-grained visual information embedded in raw observations. [6] further explores long-term memory under dynamic environments by preserving raw image observations. Nevertheless, its task formulation remains restrictive, typically evaluating an information searching system within a fixed explored space, hinders its applicability to realistic exploration scenarios. More recently, [31] focuses on extracting reusable experience from training trajectories. While effective at capturing high-level patterns, the learned experiences remain tightly coupled to previously environment, limiting their generalization ability. Motivated by these observations, we draw inspiration from human memory systems and propose to disentangle episodic and semantic memory, recalling episode to improve exploration efficiency, while following high-level decision rules in decision making.
III Methods
III-A Semantic and Physical Space for Embodied Agents
Navigation in embodied environments requires reasoning jointly over a semantic space and a physical space. Inspired by cognitive science findings suggesting structured interactions between these two spaces [21], we propose to decouple semantic memory from physical geometry, while enabling their interaction through test time retrieval.
Semantic and Physical Space. At each observation step, we extract a set of region-level semantic embeddings and a place-level image embedding using SAM [11] and CLIP [17] to form hierarchical embedding-based scene graph, following HOV-SG [23]. Global and regional embeddings are stored in a semantic memory indexed by observation time, camera pose and raw image. To interface with downstream tasks, we leverage a LLM to decompose an instruction into three semantic components: a target object, a set of relative objects, and a set of relative areas. Each component is embedded and used to retrieve the most relevant memory via similarity search over region embeddings. Retrieved memory are selected in a priority order, with target object evidence taking precedence, followed by relative objects and relative areas. While the physical space is represented by a lightweight 2D occupancy map constructed from depth observations and camera poses.
Retrieval-Time Inter-Space Interaction. Unlike prior approaches such as HOV-SG, which tightly couple semantic representations with multi-view geometric fusion, our semantic memory is updated independently of persistent geometry. Specifically, we only link semantic space to the physical space at test time by projecting the camera poses of retrieved observations onto the occupancy grid. The interaction allowing semantic concepts to be flexibly grounded without continuous online fusion, and preserves task-relevant semantic cues while keeping the physical representation compact and efficient.
III-B Human-Inspired Memory for Embodied Exploration
Embodied agents are expected to learn and adapt from their past experiences. Based on this, we propose a non-parametric method to model human-inspired memory for embodied exploration and question answering.
@@ III-B Human-Inspired Memory for Embodied Exploration
| Commercial MLLM Exploration | ||||||||||||||||
@@
Episodic Memory: Visual Reasoning for Episodic Recall. Episodic memory records episode-level experiences in a detailed and temporally grounded manner. Each episodic memory instance consists of both a semantic space and a physical space, which encode the observed scene from complementary perspectives, as described in Section III-A. As a fine-grained memory type, episodic memory preserves information at the granularity of navigation steps.
@@
Mapping-based approaches [13] explicitly align and merge memories by estimating relative transformations, which are highly sensitive to optimization failures and are prone to cascading errors during long-horizon changing and sensor noises. While we humans recall past experiences through associative and semantic similarity rather than rigid metric fusion of memories. Hence, in contrast to mapping methods, we propose a gentle episodic recall strategy supported by visual reasoning capability.
@@
Let composes the cognitive map, where represent the retrieved camera poses, and denotes the set of frontier points. For each frontier , let indicate whether it can lead to unexplored regions, and denote its spatial distance to semantically relevant landmarks. The current and previous exploration maps are jointly provided to a MLLM, outputing pruned frontiers in a visual reasoning manner:
Here, another retrieval-time interaction between physical and semantic space is expressed through frontiers pruning operation. Through this reasoning-centric approach, episodic memory informs exploration decisions without requiring explicit geometric fusion, resulting in a stable exploration strategy.
Semantic Memory: Program-Style Enhanced Rule Extraction. Semantic memory captures long-term, consolidated knowledge distilled from an agent’s past experiences, including reusable knowledge and preferred response patterns. Unlike episodic memories, which rely on fine-grained details and are inherently tied to specific environments, our semantic memory aims to extract reusable rules from dilemmas encountered during reasoning execution, with the core objective of capturing solutions that are transferable across environments. However, pioneer experiments reveal that, directly summarizing embodied agents reasoning logs are often overly entangled with environments-specific details that fail to reuse and compactly encode essential knowledge.
To address these limitations, we follow [8] and introduce a structured abstraction strategy based on pseudocode-style workflows. Specifically, we prompt an LLM to analyze the execution trace of a task by prioritizing high-level structural information. The model first identifies variables and functions that govern the workflows, and then organizes them into a pseudocode that explicitly captures control flow and decision logic.
Unlike prior work [8, 14], where workflows are typically short, embodied agent often involves long-horizon execution with extended decision sequences. In this situation, humble LLM assistant tends to over-annotate functions with learned rules that are weakly correlated with task success, obscuring the truly critical decision points. To prevent this, we introduce decision deviation detection mechanism on robots’ trajectories, following [29]. Specifically, given the robot state at time step and the corresponding ground-truth trajectory , we compute:
where measures the Euclidean distance between the agent’s current position and the ground-truth trajectory. Interactively setting a predefined threshold , we identify a decision deviation at timestep if
We randomly stop the detection at threshold when the number of decision deviation is in the range of 3-5 to form set . These deviation timesteps signal potential errors or departures from the expected behavior, allowing us to ‘debug’ the ‘program’.
@@
Finally, we can abstracts experiential knowledge into structured, high-level rules using canonical forms such as if–then, situation–suggestion and problem–solution key-value pairs, where each pair belongs to one specific variables or functions.
At the test-time, the task-similarity retrieved semantic memory is loaded into the system prompt. Details can be found in the appendix.
Meta Memory: Inspired by the success of cognitive process modeling [2], we implement meta memory via cognitive states automatic switching. Specifically, we pre-define a set of discrete cognitive states that govern high-level reasoning behavior, including: exploration, target verification, double checking, ready to answer. See Table in the appendix for concrete prompt examples associated with each cognitive state.
@@ IV-A Active Embodied Question Answering
To evaluate the effectiveness of our memory design, we conduct experiments on the Open-EQA benchmark [15, 16, 18]. Here, we compare our approach against other MLLM-based baselines on the Open-EQA benchmark, assessing both exploration efficiency and performance on answering complex open-ended questions. In addition, our semantic memory is distilled from 184 episodes in the training split, with a category distribution aligned with that of the test set.
@@
Analysis. As shown in Table I, our method consistently outperforms all baseline approaches when using the GPT-4o backbone. Specifically, the overall LLM-Match improves from 58.3% to 65.6%, while LLM-Match×SPL increases substantially from 37.3% to 48.7%, indicating both higher answer accuracy and improved exploration efficiency.
Notably, significant gains are observed across most task categories: Object recognition improves from 45.0% to 62.0%, functional reasoning from 65.0% to 66.4%, and world knowledge from 35.0% to 54.5%. On the one hand, our method achieves higher LLM-Match×SPL across all categories, demonstrating that episodic memory recall guided by visual reasoning effectively directs exploration toward more goal-relevant regions, resulting in more purposeful and efficient navigation. On the other hand, the pronounced improvements in world knowledge and functional recall highlight our semantic memory supporting the reasoning-intensive and cognitively complex tasks. Additionally, strong improvements in object recognition and object localization can be attributed to our framework’s reduced reliance on specialized perception models [5], enabling better and more flexible object identification through semantic space rather than purely textual cues.
IV-B Life-Long Visual Language Navigation
GOAT-Bench [10] is a lifelong visual navigation benchmark designed around a multi-target Go-to-AnyThing (GOAT) task. In this setting, an agent is required to navigate sequentially to multiple target objects within previously unseen environments. Targets are specified in diverse modalities, including object category names (e.g., refrigerator), natural language descriptions (e.g., toy next to the dresser), or goal images.
Following prior baseline methods [31, 27], we evaluate our approach on a -scale subset of the val-unseen split, which contains one episode per scene across 36 scenes, resulting in a total of 278 navigation subtasks. In addition, our semantic memory is constructed from the 36 training scenes.
@@ IV-B Life-Long Visual Language Navigation
| ✓ | 60.9 | 42.1 |
Metrics. We adopt Success Rate and Success weighted by Path Length (SPL) as evaluation metrics. An episode is deemed successful when the agent terminates within 1 m of the navigation target. SPL further measures efficiency by weighting successful episodes according to the length of the navigation path.
Analysis. As shown in Table II, our method outperforms most baseline methods, demonstrating its ability to effectively leverage knowledge accumulated from both episode memory and semantic memory, as evidenced by consistently higher Success Rate and SPL scores. In particular, when compared with 3D-Mem under the GPT-4o backbone, our method achieves a notable improvement in life-long navigation, with the Success Rate increasing from 65.1% to 72.8%, and SPL improving from 49.3% to 56.1%, indicating more accurate and efficient goal-directed exploration. Under the Qwen backbone, our method slightly underperforms ReEXplore in SPL metric. We attribute this primarily to the relatively weaker visual reasoning capability of Qwen backbone when operating on semantic-labeled exploration maps, which limits the effectiveness of frontier pruning operation during exploration.

### v2

# HIMM: Human-Inspired Long-Term Memory Modeling for Embodied Exploration and Question Answering

@@ Abstract
Deploying Multimodal Large Language Models as the brain of embodied agents poses challenging when reasoning long-horizon observations. Existing memory-augmented approaches typically compress observations into textual summaries losing fine-grained information, and overlook the inherent differences across memory types. To address these limitations, we propose HIMM, a memory framework that explicitly models episodic and semantic memory for embodied exploration and question answering. Our approach recalls episodic experiences based on semantic similarity and verifies them through visual reasoning on exploration maps, enabling efficient reuse of past observations without strict geometric alignment. Meanwhile, we introduce a program-style rule extraction mechanism that transforms experiences into structured, reusable semantic memory, facilitating cross-environment generalization. Our methods outperforms prior methods on on embodied exploration and question answering benchmarks, yielding a 7.3% LLM-Match gain and an 11.4% LLM-MatchSPL gain on A-EQA, as well as +7.7% success rate and +6.8% SPL on GOAT-Bench. Analyses reveal that our episodic memory primarily improves exploration efficiency, while semantic memory strengthens complex reasoning of embodied agents.
@@ I Introduction
In recent years, Multimodal Large Language Models (MLLMs) have advanced rapidly, demonstrating strong reasoning and perception capabilities across a wide range of domains. Despite this progress, deploying MLLMs as the core module of embodied agents in real-world environments remains highly challenging. A key difficulty lies in enabling agents to efficiently process long-horizon observations while retaining task-relevant information under limited context lengths. To address this issue, recent works introduce various forms of working memory, allowing agents to retrieve a small subset of past observations instead of conditioning on the entire sensory history [25, 27, 29]. By coupling task-relevant information retrieval with frontier-based exploration strategies, these methods achieve efficient and effective performance in embodied exploration and question answering.
@@
However, most existing approaches remain confined to information within the current episode, limiting their ability to leverage accumulated experience over time. In contrast, real-world embodied agents are expected to continuously integrate observations across episodes and exploit long-term knowledge acquired from past interactions. Approaches such as [32, 33] attempt to address this by summarizing cross-episode information into hierarchical textual descriptions that guide future exploration. Meanwhile, [5] introduces cross-episode memory into the reasoning loop. Nevertheless, its task formulation remains restrictive, typically evaluating an MLLM-based information-collection system within a fixed explored space derived from a previous timeline. Despite these efforts, existing memory architectures remain fundamentally rigid, and they suffer from two major limitations. First, this design inevitably discards a substantial amount of spatial and visual cues in raw observations, limiting the agent’s memory capacity. Second, most methods adopt a single memory representation to model heterogeneous memory systems, which fails to capture the distinct characteristics and functions of different memory types, as extensively studied in cognitive psychology [22].
@@
These limitations become even more pronounced when moving from simulation to realistic dynamic environments, where two additional fundamental challenges arise. First, fusion maps across time, whether represented as persistent nodes [27] or voxel grids [24], becomes inherently difficult in the presence of environmental changes and noise. Second, information retrieved from past does not necessarily reflect the current state of the environment. Based on these observations, we argue that embodied agent should neither rely on rigid long-term geometric fusion nor blindly trust past episodic information. Instead, it should recall prior experiences in a soft and associative manner while continuously grounding them in active exploration and perception, closely mirroring how humans reason about and interact with dynamic environments.
Motivated by these insights, we propose HIMM, a non-parametric memory framework inspired by human memory systems, which explicitly model episodic and semantic memory for embodied exploration and question answering. We propose a retrieval-first and reasoning-assisted paradigm. Retrieved memory snapshot are selectively grounded and verified through visual reasoning, ensuring reliable reuse in long-horizon timeline. Complementarily, in contrast to leveraging environment-specific and free-form textual summaries, our method distills structured, high-level, and reusable semantic memory through pseudocode-like decision workflows and decision deviations detection. This dual-system design enables stable long-term knowledge accumulation while supporting robust exploration and long-horizon reasoning. We summarize our contributions as follows:
We introduce HIMM, a memory framework that explicitly models episodic and semantic memory for embodied agents, enabling self-evolving without additional training.
We propose to ground retrieved episodic memories through visual reasoning over memory snapshots and exploration maps, together with a program-style rule extraction mechanism that enhances complex embodied reasoning.
@@ II-A MLLM-based Embodied Agents
Despite these advances, foundation models still remain computationally expensive, and naively processing uniformly sampled frames is often suboptimal for questions that require selectively identifying informative observations for exploration and reasoning. To address these challenges, [6, 25, 27, 35] introduces CLIP [16] and SAM-based [10] open-vocabulary scene graphs to structure semantics for MLLM-driven exploration. While [29] further incorporates raw images to enrich memory, its key-frame selection remains constrained by detectors and lacks explicit geometric modeling, leading to inefficient long-horizon exploration. In contrast to existing geometric-semantic graph methods often depend on heavy multi-frame fusion, our approach marks a shift toward decoupling semantic and geometric spaces, enabling retrieval-time interaction that supports adaptive cross-episode recalling.
@@ II-B Experience and Memory for Embodied Agents
Research on memory mechanisms for embodied agents seeks to enable the storage, retention, and retrieval of past experiences, facilitating the transition from purely reactive systems to agents capable of maintaining context and autonomous adaptation. [1] and [32] adopt descriptive text, temporal and spatial cues to manage continuously growing histories, allowing robots to recall previously visited regions and reduce redundant exploration. However, memory built primarily on image and video captioning discard a substantial amount of fine-grained visual information embedded in raw observations. [5] further explores long-term memory under dynamic environments by preserving raw image observations. Nevertheless, its task formulation remains restrictive, typically evaluating an information searching system within a fixed explored space, hinders its applicability to realistic exploration scenarios. More recently, [33] focuses on extracting reusable experience from training trajectories. While effective at capturing high-level patterns, the distilled experiences remain tightly coupled to previously environment, limiting their generalization ability and reasoning enhancement. Motivated by these observations, we draw inspiration from human memory systems and propose to disentangle episodic and semantic memory, recalling episode to improve exploration efficiency, while following high-level decision rules in decision making.
@@
| Commercial MLLM-based Exploration | ||||||||||||||||
@@
III Methods
Our overall pipeline is illustrated in Fig. 1. Given a posed RGB-D observation, our framework constructs semantic and physical space to represent the scene (Section III-A), Next, episodic memory is recalled by visual reasoning to guide exploration, while program-sytle extracted semantic memory is fed into LLM to improve rational (Section III-B).
III-A Semantic and Physical Space for Embodied Agents
Navigation in embodied environments requires reasoning jointly over a semantic space and a physical space. Inspired by cognitive science findings suggesting structured interactions between these two spaces [23], we propose to decouple semantic memory from physical geometry, while enabling their interaction through test-time retrieval.
Semantic and Physical Space. At each observation step, a set of region-level semantic embeddings and a place-level image embedding are extracted using SAM and CLIP to form hierarchical embedding-based scene graph, following HOV-SG [25]. Global embedding, regional embeddings, region 3D-boxes and raw image are stored in a semantic space indexed by camera pose, as depicted in Fig. 2. To interface with downstream tasks, we leverage a LLM to decompose an instruction into three goal components: a target object, a set of relative objects, and a set of relative areas. Each component is embedded and used to query region embeddings in the semantic space. Retrieved embeddings are ranked in a priority order, with target object evidence taking precedence, followed by relative objects and relative areas. While the physical space is represented by a lightweight 2D occupied exploration map constructed from depth observations and camera poses.
Retrieval-Time Inter-Space Interaction. In constrast to prior approaches such as [25], which tightly couple semantic representations with multi-view geometric fusion, our semantic memory is updated independently of persistent geometry. Specifically, we only link semantic space to the physical space by rendering the camera poses of retrieved observations onto the occupancy grid. The interaction allowing semantic concepts to be flexibly grounded without continuous online fusion, and preserves task-relevant semantic cues while keeping the physical representation compact and efficient.
III-B Human-Inspired Memory for Embodied Exploration
Embodied agents are expected to learn and adapt from their past experiences. Episodic and semantic memory play essential roles in human cognition [22]. This inspires that instead of using a single memory representation to model heterogeneous memory systems, it is important to design different construction and retrieval strategies for each type of memory to better capture their distinct functions.
Episodic Memory: Visual Reasoning for Episodic Recall. Episodic memory records episode-level experiences in a detailed and temporally grounded manner [22]. Each episodic memory instance consists of both a semantic space and a physical space, which encode the observed scene from complementary perspectives, as described in Section III-A. As a fine-grained memory type, episodic memory preserves information at the granularity of navigation steps.
@@ III-B Human-Inspired Memory for Embodied Exploration
Mapping-based approaches [12] explicitly align and merge memories by transformations estimation, which are highly sensitive to alignment failures and are prone to cascading errors during long-horizon changing and noises. While we humans recall past experiences through associative and semantic similarity rather than rigid metric fusion of memories. Hence, in contrast to these methods, we propose a gentle episodic recall strategy supported by visual reasoning capability.
@@
Let composes the exploration map, where represent the retrieved camera poses, and denotes the set of frontier points. For each frontier , let indicate whether it can lead to unexplored regions, and denote its spatial distance to semantically relevant landmarks. The current and previous exploration maps are jointly provided to a MLLM, outputing pruned frontiers in a visual reasoning manner:
Here, another retrieval-time interaction between semantic and physical space is expressed through frontiers pruning operation. Through this reasoning-centric approach, episodic memory informs exploration decisions without requiring explicit geometric fusion, resulting in a stable exploration strategy.
Semantic Memory: Program-Style Enhanced Rule Extraction. Semantic memory captures long-term, consolidated knowledge distilled from an agent’s past experiences, including reusable knowledge and preferred response patterns [21]. Unlike episodic memories, which rely on fine-grained details and are inherently tied to specific environments, semantic memory aims to extract reusable rules from dilemmas encountered during reasoning execution, with the core objective of capturing solutions that are transferable across environments. However, previous experiments [13] reveal that, directly summarizing agents’ reasoning logs are often overly entangled with details that fail to reuse and compress essential knowledge.
To address these limitations, we draw inspiration from [7] and introduce a program-style enhanced rule extraction strategy. Specifically, we prompt an LLM to analyze the execution pseudocode of a task to prioritize high-level structural information. The model first identifies variables and functions that govern the workflows, and then organizes them into a pseudocode that explicitly captures control flow and decision logic. In constract to [7, 13], where workflows are typically short, embodied agent often involves long-horizon execution with extended decision sequences. In this situation, LLMs tend to over-annotate functions with learned rules that are weakly correlated with task success, obscuring the truly critical decision points. To prevent this, we introduce decision deviations detection mechanism on robots’ trajectories [31]. Specifically, given the robot state at timestep and the corresponding ground-truth trajectory , we compute:
where measures the Euclidean distance between the agent’s current position and the ground-truth trajectory from training set. In constract to previous work using predefined threshold, we interactively set threshold to identify decision deviations at timestep if
Then, we randomly stop the detection at threshold when the number of decision deviation is in the range of 3-5 to form set . These deviation timesteps signal potential errors or departures from the expected behavior, allowing embodied agent to ‘debug’ the ‘program’.
@@
Finally, experience semantic knowledge can be abstracted into structured, high-level rules using canonical forms such as if–then, situation–suggestion and problem–solution key-value pairs, where each pair belongs to one specific variables or functions.
At the test-time, the question-similarity retrieved semantic memory is loaded into the system prompt to constrain decision-making process. Specifically, MiniLM-L6-v2 variant from Sentence-Transformer [18] is used to extract embeddings for querying question in semantic memory, and compute cosine similarity to retrieve the top-K.
Meta Memory: We leverage the success of cognitive process modeling [2, 5] and introduce meta memory via cognitive states automatic switching to build the foundation of embodied agent. Specifically, pre-defined discrete cognitive states include: exploration, target verification, target approaching, check ready to answer.
@@ IV-A Active Embodied Question Answering
To evaluate the effectiveness of our memory design, we conduct experiments on the A-EQA benchmark [14] which built on habitat [15, 17]. Here, we compare our approach against other MLLM-based baselines on the A-EQA benchmark, assessing both exploration efficiency and performance on answering complex open-ended questions. In addition, our semantic memory is distilled from 184 episodes in the training split, with a category distribution aligned with that of the test set.
@@
Analysis. Table I shows that our method surpasses all baselines with the GPT-4o backbone. Specifically, the overall LLM-Match improves from 58.3% to 65.6%, and LLM-Match×SPL increases substantially from 37.3% to 48.7%, indicating both higher answer accuracy and improved exploration efficiency.
@@
| ✓ | 60.9 | 41.1 |
Notably, significant gains are observed across most task categories: Object recognition improves from 45.0% to 62.0%, functional reasoning from 65.0% to 66.4%, and world knowledge from 35.0% to 54.5%. On the one hand, our method achieves higher LLM-Match×SPL across all categories. As further illustrated in Fig. 4, episodic memory recall guided by visual reasoning effectively steers exploration toward goal-relevant regions, resulting in more purposeful navigation. On the other hand, the pronounced improvements in world knowledge and functional recall highlight our semantic memory supporting the reasoning-intensive and cognitively complex tasks. Fig. 5 illustrates how semantic memory constrains the agent’s reasoning during question answering. Task-relevant semantic memory encapsulates high-level prior experience, reducing the arbitrariness of MLLM reasoning and constraining the agent to adopt more appropriate strategies, thereby improving answering performance.
IV-B Life-Long Visual Language Navigation
GOAT-Bench [9] is a lifelong visual navigation benchmark designed around a multi-target navigation task. In this setting, an agent is required to navigate sequentially to multiple target objects within previously unseen environments. Targets are specified in diverse modalities, including object category names (e.g., refrigerator), natural language descriptions (e.g., toy next to the dresser), or goal images. Following prior baseline methods [33, 29], we evaluate our approach on a -scale subset of the val-unseen split, which contains one episode per scene across 36 scenes, resulting in a total of 278 navigation subtasks. In addition, our semantic memory is constructed on the 36 training scenes.
Metrics. Success Rate and Success weighted by Path Length (SPL) is adpoted as evaluation metrics. An episode is deemed successful when the agent terminates within 1 m of the navigation target. Similar to OpenEQA, SPL measures efficiency by weighting successful episodes according to the length of the navigation path.
Analysis. As shown in Table II, our method outperforms most baseline methods, demonstrating its ability to effectively leverage knowledge accumulated from both episode memory and semantic memory, as evidenced by consistently higher Success Rate and SPL scores. Compared with 3D-Mem under the GPT-4o backbone, our method achieves a notable improvement in life-long navigation. The Success Rate increases from 68.9% to 72.8%, and SPL rises from 48.9% to 56.1%, indicating more accurate and efficient goal-directed exploration. Under the Qwen backbone, our method slightly underperforms ReEXplore in SPL metric. We attribute this primarily to the relatively weaker visual reasoning capability of Qwen backbone when operating on semantic-labeled exploration maps, which limits the correctness of frontier pruning.